# Hybrid Inflation, Reheating and Dark Radiation in a IIB perturbative moduli stabilization scenario

### v3

@@
[…]
${}^{a}$ Center for Fundamental Physics and School of Mathematics and Physics, Hubei Polytechnic University
@@
${}^{b}$ Physics Department, University of Ioannina
@@ 3 The effective potential
Figure 1: Plots of the potential along the volume direction. The left panel shows the F-term potential, while in the right panel, the D-term potential has also been included. We choose $\xi_{0}=10$, $\eta_{0}=-0.92$, $S=0$, $\varphi_{1,o}=\varphi_{2,o}=M$, $\kappa=0.1$ and $\gamma=1.$ Here $x$ represents the volume, $x\equiv\mathcal{V}$.
@@
$V_{D}=\sum_{i=1}^{3}\frac{g_{D7_{i}}^{2}}{2}\left(Q_{i}\partial_{T_{i}}K+\sum_{j}q_{j}|\Theta^{j}|^{2}\right)^{2},$ (3.5)
where ${g_{D7_{i}}}=({\rm Re}T_{i})^{-1}$ and $Q_{j},q_{j}$ are “charges”, while $\Theta^{j}$ represent matter fields charged under the $U(1)$ gauge factors. Some aspects regarding strings at $D7$ intersections in particular have been recently discussed in [17]. The fields $\Theta^{j}$ carry charges under the $U(1)$ factors associated with the D-branes. A compelling question then arises regarding the possible contributions of these fields in the D-term, and hence, in the scalar potential. One possible scenario is that non-zero field vevs are chosen so that the D-term vanishes, and in this case, the uplift should be realized by a suitable modification (see footnote 10), or the standard procedure of introducing $\overline{D3}$ branes (see [32]). Here we make a simplified assumption that these singlets have vanishing vevs. Whether the vevs of these fields are zero or not, depends on the specific details of the effective model. One possibility is to assume that in the effective field theory limit, all these fields $\Theta^{j}$ are minimized at $\langle\Theta^{j}\rangle=0$. Alternatively, even in the case of non-zero vevs, however, it might happen that there could be accidental cancellations that can reduce the significance of their contributions. Some related discussions regarding these issues can be found also in [33]. Here, since the D-term potential is only used to uplift the non-supersymmetric AdS vacuum of the F-term potential, following closely previous works (see for example [10], [16] and [31]) we make the reasonable assumption that moduli fields dominate over other fields vevs. Therefore, it is adequate for our purposes to assume that the flux induced D-term piece $\propto Q_{i}\partial_{T_{i}}K$ dominates, and therefore we minimize the potential by setting $\langle\Theta^{j}\rangle$ vevs zero. For $\langle\Theta^{j}\rangle=0$ the second term in the parenthesis vanishes and each component of the D-term acquires a simple -model independent- form $V_{D_{i}}\approx Q_{i}^{2}/\tau_{i}^{3}$. Then, the total D-term potential, being the sum of three components, is approximated by [10]
@@
Here, $d_{i}$ are positive constants related to the charges $d_{i}\sim Q_{i}^{2}>0$. Furthermore, using the volume formula ${\cal V}^{2}=\tau_{1}\tau_{2}\tau_{3}$, we have substituted the modulus $\tau_{2}$ with its equivalent $\tau_{2}={\cal V}^{2}/(\tau_{1}\tau_{3})$ Then, the total effective potential is the sum $V_{\rm eff}=V_{F}+V_{D}$ which can be minimized with respect to $\mathcal{V}$ and the two remaining Kähler moduli $\tau_{1,2}$. We have assumed that the F-part of the potential depends on the total volume, hence the explicit dependence of $V_{\rm eff}$ on $\tau_{1},\tau_{3}$ only comes through the $V_{D}$ part. It is found that the two minimization conditions with respect to $\tau_{1,3}$ determine the ratios between the moduli, i.e., $\left(\frac{\tau_{i}}{\tau_{j}}\right)^{3}=\frac{d_{i}}{d_{j}}$ [34]. Expressed in terms of the stabilized total volume $\mathcal{V}$, the conditions for the two $\tau_{i}$ can be written as
@@
By choosing $\varphi_{1}=\varphi_{2}=\varphi$ and $\alpha=\beta$, the D-term contribution from the Higgs fields vanishes. Then the effective potential has the following form
@@ 3.1 Inflationary phase
is present at tree level, thus SUSY is broken during inflation. When SUSY breaks, splitting between fermionic and bosonic mass multiplets is created and contributions to radiative corrections occur. Following [35, 36], the soft terms are [^11] Note that we use S for both the canonically normalized field and the non-canonically one. The scaling factor required for the canonically normalized field is adjusted by multiplying ${\cal{V}}_{0}^{2/3}/(3a\gamma)$.
@@
It should be observed that the Coleman-Weinberg correction is computed along the inflationary trajectory, wherein the field $S$ takes a non-zero value while the field $\varphi$ remains fixed at zero. Furthermore, we set $V_{o}=3.2\times 10^{4}$ for the volume field, and the mass spectrum solely depends on $S$. The detailed calculations are provided in Appendix A.
@@
where prime denotes the derivative with respect to $y$. In the slow-roll approximation, the scalar spectral index $n_{s}$, the tensor-to-scalar ratio $r$ and the running of the scalar spectral index $\alpha_{s}\equiv dn_{s}/d\ln k$ are given by,
@@
Figure 3: Variations of $r$ and $M$ in $\kappa-m_{3/2}$ plane. The upper line has $M$ fixed at a value equal to $M_{string}$. The other two lines correspond to lower $M$ values, as indicated in the plot.
@@ 4 Reheating and dark radiation
$K_{i\bar{j}}=\frac{1}{4}{\rm diag}\left(\frac{1}{\tau_{1}^{2}},\frac{1}{\tau_{3}^{2}},\frac{1}{\tau_{3}^{2}}\right).$ (4.3)
Therefore, Eq (4.1) can be rewritten as,
@@
$u=\frac{u_{1}-u_{2}}{\sqrt{2}}=\frac{1}{2}\log{\frac{\tau_{1}}{\tau_{2}}},\;v=\frac{u_{1}+u_{2}-2u_{3}}{\sqrt{6}}=\frac{1}{\sqrt{3}}\log{\frac{\tau_{1}\tau_{2}}{\tau_{3}^{2}}}\;.$
We reverse the relations
$\left(\begin{array}[]{c}u_{1}\\ u_{2}\\ u_{3}\\ \end{array}\right)=\left(\begin{array}[]{ccc}\frac{1}{\sqrt{3}}&\frac{1}{\sqrt{2}}&\frac{1}{\sqrt{6}}\\ \frac{1}{\sqrt{3}}&-\frac{1}{\sqrt{2}}&\frac{1}{\sqrt{6}}\\ \frac{1}{\sqrt{3}}&0&-\sqrt{\frac{2}{3}}\\ \end{array}\right)\left(\begin{array}[]{c}t\\ u\\ v\\ \end{array}\right)$
while substituting to the Lagrangian part for axions we have
${\cal L}$ $\supset$ $\frac{1}{\sqrt{3}}t\left(\partial_{\mu}a_{1}\partial^{\mu}a_{1}+\partial_{\mu}a_{2}\partial^{\mu}a_{2}+\partial_{\mu}a_{3}\partial^{\mu}a_{3}\right)$ (4.4)
$+\frac{1}{\sqrt{2}}u\left(\partial_{\mu}a_{1}\partial^{\mu}a_{1}-\partial_{\mu}a_{2}\partial^{\mu}a_{2}\right)$
$+\frac{1}{\sqrt{6}}v\left(\partial_{\mu}a_{1}\partial^{\mu}a_{1}+\partial_{\mu}a_{2}\partial^{\mu}a_{2}-2\partial_{\mu}a_{3}\partial^{\mu}a_{3}\right).$
@@
where $m_{u}$ is the modulus mass. In the large volume scenario a distinct hierarchy of mass scales is generated, (for details see Ref [40]). The mass eigenstates after diagonalization in units of $M_{p}=1$ can be written as,
@@
where $g_{*}(T_{R}h)$ is the effective degree of freedom at the time of reheating the Universe. The measured values of $N_{\text{eff}}$ require $\Delta N_{\text{eff}}\lesssim 0.95$ at the $95\%$ confidence level, which translates into a bound on the parameters of the model, $a\cdot\lambda\gtrsim 0.1688$. In Fig. 3, the parametric space predicted by the inflationary analysis has $0.86\leq a\leq 1$. This range aligns with the values predicted by $\Delta N_{\text{eff}}\lesssim 0.95$ and is consistent with dark radiation. Within this parametric space, we obtain a tensor-to-scalar ratio values of $r\leq 1.4\times 10^{-4}$, $1\times 10^{15}\lesssim M\lesssim 1.3\times 10^{16}$ GeV, and a soft mass parameter $4.6\times 10^{5}\lesssim{\color[rgb]{0,0,0}M_{S_{c}}}\lesssim 1\times 10^{11}$ GeV.
@@ 5 Conclusions
Appendix A Coleman-Weinberg Corrections
@@
$\small V_{\text{eff}}=\frac{\kappa^{2}\alpha\mid\left(M^{2}-\Phi_{1}\Phi_{2}\right)\mid^{2}+2\gamma\kappa^{2}S^{2}(\alpha\Phi_{1}\Phi_{1}^{\dagger}+\beta\Phi_{2}\Phi_{2}^{\dagger})}{3a\alpha\beta\gamma\mathcal{V}^{4/3}}+\frac{3W_{0}^{2}(2\eta_{0}\log(\mathcal{V})-8\eta_{0}+\xi_{0})}{2\mathcal{V}^{3}}+\frac{d}{\mathcal{V}^{2}},$ (A.1)
@@
$M_{S}^{2}=\begin{pmatrix}\frac{\kappa^{2}S^{2}}{6a\beta\mathcal{V}^{4/3}}&0&-\frac{\kappa^{2}M^{2}}{3a\gamma\mathcal{V}^{4/3}}&0&0&0&0\\ 0&\frac{\kappa^{2}S^{2}}{6a\beta\mathcal{V}^{4/3}}&0&\frac{\kappa^{2}M^{2}}{3a\gamma\mathcal{V}^{4/3}}&0&0&0\\ -\frac{\kappa^{2}M^{2}}{3a\gamma\mathcal{V}^{4/3}}&0&\frac{\kappa^{2}S^{2}}{6a\alpha\mathcal{V}^{4/3}}&0&0&0&0\\ 0&\frac{\kappa^{2}M^{2}}{3a\gamma\mathcal{V}^{4/3}}&0&\frac{\kappa^{2}S^{2}}{6a\alpha\mathcal{V}^{4/3}}&0&0&0\\ 0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0\\ 0&0&0&0&0&0&\frac{28\kappa^{2}M^{4}}{27a\gamma\mathcal{V}^{\frac{10}{3}}}+\cdots\\ \end{pmatrix}$ (A.2)
@@
$\small M_{S}^{2}=\left[0,0,\frac{\kappa^{2}M^{2}\left(y^{2}-\frac{2\alpha}{\gamma}\right)}{6a\mathcal{V}^{4/3}\alpha},\frac{\kappa^{2}M^{2}\left(y^{2}-\frac{2\alpha}{\gamma}\right)}{6a\mathcal{V}^{4/3}\alpha},\frac{\kappa^{2}M^{2}\left(y^{2}+\frac{2\alpha}{\gamma}\right)}{6a\mathcal{V}^{4/3}\alpha},\frac{\kappa^{2}M^{2}\left(y^{2}+\frac{2\alpha}{\gamma}\right)}{6a\mathcal{V}^{4/3}\alpha},\quad\frac{28\kappa^{2}M^{4}}{27a\gamma\mathcal{V}^{10/3}}+\cdots\right]$ (A.3)
@@
$M_{F_{ij}}=\begin{pmatrix}-\frac{6a\gamma\kappa M^{2}S}{\mathcal{V}^{2/3}}+\frac{9a^{2}\gamma^{2}S^{2}}{\mathcal{V}^{4/3}}+\cdots&0&0&\frac{2kM^{2}}{\mathcal{V}}+\frac{2a\gamma S\left(kM^{2}S-4\right)}{\mathcal{V}^{5/3}}+\cdots\\ 0&0&\kappa S&0\\ 0&\kappa S&0&0\\ \frac{2\kappa M^{2}}{\mathcal{V}}+\frac{2a\gamma S\left(\kappa M^{2}S-4\right)}{\mathcal{V}^{5/3}}+\cdots&0&0&\frac{6}{\mathcal{V}^{2}}+\frac{34}{3}a\gamma S^{2}\left(\frac{1}{\mathcal{V}}\right)^{8/3}+\cdots\\ \end{pmatrix}.$ (A.5)
@@
$M_{F_{ij}}^{2}=\left[\kappa^{2}M^{2}y^{2},\kappa^{2}M^{2}y^{2},\frac{36a^{2}\kappa^{2}M^{6}y^{2}\gamma^{2}}{\mathcal{V}^{4/3}}+\cdots,\frac{36a^{2}\kappa^{2}M^{6}y^{2}\gamma^{2}}{\mathcal{V}^{4/3}}+\cdots\right].$ (A.6)
The fermionic mass $\frac{36a^{2}\kappa^{2}M^{6}y^{2}\gamma^{2}}{\mathcal{V}^{4/3}}$ and the bosonic mass $\,\frac{28\kappa^{2}M^{4}}{27a\gamma\mathcal{V}^{10/3}}$ are small compared to other masses, so their contribution to the Coleman-Weinberg potential is suppressed. Note that we use the symbol $y$ for both the canonically, and the non-canonically normalized field. The scaling factor required for the canonically normalized field is adjusted by multiplying ${\cal{V}}_{0}^{2/3}/(3a\gamma)$. Additionally, we focus on the region where $(y^{2}\gg 2\alpha/\gamma)$. The effective contribution of the one-loop radiative corrections can be calculated using the Coleman-Weinberg formula
$\Delta V_{\text{1-loop}}=\frac{1}{64\pi^{2}}\left[M_{S}^{4}\log\left(\frac{M_{S}^{2}}{Q^{2}}\right)-2M_{F}^{4}\log\left(\frac{M_{F}^{2}}{Q^{2}}\right)-\frac{3}{2}\left(M_{S}^{4}-2M_{F}^{4}\right)\right]$ (A.7)
$\Delta V_{\text{1-loop}}=\frac{\kappa^{4}M^{4}y^{4}}{144\pi^{2}a^{2}\gamma^{2}{\cal V}_{o}^{4/3}}\left[F(y)-\left(\frac{1}{54a^{2}\alpha^{2}\gamma^{2}}-\frac{3}{2}{\cal V}_{o}^{8/3}\right)\right],$ (A.8)
@@
$F(y)=\frac{1}{81a^{2}\alpha^{2}\gamma^{2}}\log\left(\frac{\kappa^{2}M^{2}y^{2}}{27a^{2}\alpha\gamma^{2}Q^{2}{\cal V}_{o}^{2/3}}\right)-{\cal V}_{o}^{8/3}\log\left(\frac{\kappa^{2}M^{2}y^{2}{\cal V}_{o}^{2/3}}{3a\gamma Q^{2}}\right)~{}.$ (A.9)

### v4

@@
Waqas Ahmeda  [^1] E-mail: waqasmit@hbpu.edu.cn, Athanasios Karozasb  [^2] E-mail: athanasioskarozas@gmail.com, George K. Leontarisb  [^3] E-mail: leonta@uoi.gr, Ilias Tavellarisb  [^4] E-mail: i.tavellaris@uoi.gr
a Center for Fundamental Physics and School of Mathematics and Physics, Hubei Polytechnic University
@@
b Physics Department, University of Ioannina
@@ 3 The effective potential
Figure 1: Plots of the potential along the volume direction. The left panel shows the F-term potential, while the right panel also includes the D-term potential. We choose $\xi_{0}=10$, $\eta_{0}=-0.92$, $S=0$, $\varphi_{1,o}=\varphi_{2,o}=M$, $\kappa=0.1$ and $\gamma=1.$ Here $x$ represents the volume, $x\equiv\mathcal{V}$.
@@
${\color[rgb]{0,0,0}V_{D}=\sum_{i=1}^{3}\frac{g_{D7_{i}}^{2}}{2}\left(Q_{j}\partial_{T_{j}}K+\sum_{j\neq i}q^{j}_{i}|\Theta^{j}_{i}|^{2}\right)^{2},}$ (3.5)
where ${g^{2}_{D7_{i}}}=({\rm Re}T_{i})^{-1}$ and $Q_{j},q^{j}_{i}$ are “charges”, while $\Theta^{j}_{i}$ represent matter fields charged under the $U(1)$ gauge factors. Some aspects regarding strings at $D7$ intersections in particular have been recently discussed in [17]. The fields $\Theta^{j}_{i}$ carry charges under the $U(1)$ factors associated with the D-branes. A compelling question then arises regarding the possible contributions of these fields in the D-term, and hence, in the scalar potential. One possible scenario is that non-zero field vevs are chosen so that the D-term vanishes, and in this case, the uplift should be realized by a suitable modification (see footnote 10), or the standard procedure of introducing $\overline{D3}$ branes (see [32]). Here we make a simplified assumption that these singlets have vanishing vevs. Whether the vevs of these fields are zero or not, depends on the specific details of the effective model. One possibility is to assume that in the effective field theory limit, all these fields $\Theta^{j}_{i}$ are minimized at $\langle\Theta^{j}_{i}\rangle=0$. Alternatively, even in the case of non-zero vevs, however, it might happen that there could be accidental cancellations that can reduce the significance of their contributions. Some related discussions regarding these issues can also be found in [33]. Here, since the D-term potential is only used to uplift the non-supersymmetric AdS vacuum of the F-term potential, following closely previous works (see for example [10], [16] and [31]) we make the reasonable assumption that moduli fields dominate over other fields vevs. Therefore, it is adequate for our purposes to assume that the flux induced D-term piece $\propto Q_{j}\partial_{T_{j}}K$ dominates, and therefore we minimize the potential by setting $\langle\Theta^{j}_{i}\rangle$ vevs zero. For $\langle\Theta^{j}_{i}\rangle=0$ the second term in the parenthesis vanishes and each component of the D-term acquires a simple -model independent- form $V_{D_{i}}\approx Q_{i}^{2}/\tau_{i}^{3}$. Then, the total D-term potential, being the sum of three components, is approximated by [10]
@@
Here, $d_{i}$ are positive constants related to the charges $d_{i}\sim Q_{i}^{2}>0$. In our computations, we focused solely on the simplest scenario involving diagonal terms. However, in more general cases, cross-terms are also present. We observed that these cross terms exhibit the same order of magnitude in relation to the expression of $\tau_{i}$’s. Numerical analysis reveals the existence of minima in dS, albeit with slightly different coefficient values, $d_{i}$. Furthermore, using the volume formula ${\cal V}^{2}=\tau_{1}\tau_{2}\tau_{3}$, we have substituted the modulus $\tau_{2}$ with its equivalent $\tau_{2}={\cal V}^{2}/(\tau_{1}\tau_{3})$ Then, the total effective potential is the sum $V_{\rm eff}=V_{F}+V_{D}$ which can be minimized with respect to $\mathcal{V}$ and the two remaining Kähler moduli $\tau_{1,2}$. We have assumed that the F-part of the potential depends on the total volume, hence the explicit dependence of $V_{\rm eff}$ on $\tau_{1},\tau_{3}$ only comes through the $V_{D}$ part. It is found that the two minimization conditions with respect to $\tau_{1,3}$ determine the ratios between the moduli, i.e., $\left(\frac{\tau_{i}}{\tau_{j}}\right)^{3}=\frac{d_{i}}{d_{j}}$ [34]. Expressed in terms of the stabilized total volume $\mathcal{V}$, the conditions for the two $\tau_{i}$ can be written as
@@
At tree-level, the potential exhibits two flat directions when the variable S equals zero. In this scenario, one can choose different values for the variables $\phi_{1}$ and $\phi_{2}$ to create distinct minima in the potential. However, it’s essential to note that this principle applies solely when considering the F-term potential. The total potential at tree level comprises both F-term and D-term components. The D-term potential encompasses two parts: one stemming from the moduli and the other from the matter fields. To nullify the contribution from the matter fields in the D-term potential, we utilize D-flat directions where $\varphi_{1}=\varphi_{2}=\varphi$ and $\alpha=\beta$. Then the effective potential has the following form
@@ 3.1 Inflationary phase
is present at tree level, thus SUSY is broken during inflation. When SUSY breaks, splitting between fermionic and bosonic mass multiplets is created and contributions to radiative corrections occur. Following [35, 36], the soft terms are [^11] Note that the factor ${\cal{V}}_{0}^{2/3}/(3a\gamma)$ originates from the transition to the canonically normalized field $s=S\sqrt{3a\gamma}/V_{0}^{1/3}$. See also Appendix A.
@@
It should be observed that the Coleman-Weinberg correction is computed along the inflationary trajectory, wherein the field $s$ takes a non-zero value while the field $\varphi$ remains fixed at zero. Furthermore, we set $V_{o}=3.2\times 10^{4}$ for the volume field [^13] For a discussion on multifield inflation scenario in this type of models see [39]., and the mass spectrum solely depends on $s$. The detailed calculations are provided in Appendix B.
@@
where prime denotes the derivative with respect to $y$. Note that, the presence of the mass parameter $M$ in (3.20) comes from the definition of the field $y$ as the ratio $y=s/M$. In the slow-roll approximation, the scalar spectral index $n_{s}$, the tensor-to-scalar ratio $r$ and the running of the scalar spectral index $\alpha_{s}\equiv dn_{s}/d\ln k$ are given by,
@@
Figure 3: Variations of $r$ and $M$ in $\kappa-M_{s_{c}}$ plane. The upper line has $M$ fixed at a value equal to $M_{string}$. The other two lines correspond to lower $M$ values, as indicated in the plot.
@@ 4 Reheating and dark radiation
$K_{i\bar{j}}=\frac{1}{4}{\rm diag}\left(\frac{1}{\tau_{1}^{2}},\frac{1}{\tau_{2}^{2}},\frac{1}{\tau_{3}^{2}}\right).$ (4.3)
Therefore, Eq (4.1) can be rewritten as [41],
@@
$u=\frac{u_{1}-u_{2}}{\sqrt{2}}=\frac{1}{2}\log{\frac{\tau_{1}}{\tau_{2}}},\;v=\frac{u_{1}+u_{2}-2u_{3}}{\sqrt{6}}=\frac{1}{\sqrt{12}}\log{\frac{\tau_{1}\tau_{2}}{\tau_{3}^{2}}}\;.$
Finally, reversing the above relations and expressed them in matrix form we have :
$\left(\begin{array}[]{c}u_{1}\\ u_{2}\\ u_{3}\\ \end{array}\right)=\left(\begin{array}[]{ccc}\frac{1}{\sqrt{3}}&\frac{1}{\sqrt{2}}&\frac{1}{\sqrt{6}}\\ \frac{1}{\sqrt{3}}&-\frac{1}{\sqrt{2}}&\frac{1}{\sqrt{6}}\\ \frac{1}{\sqrt{3}}&0&-\sqrt{\frac{2}{3}}\\ \end{array}\right)\left(\begin{array}[]{c}t\\ u\\ v\\ \end{array}\right)\;.$
Regarding the Lagrangian part for the axions we have [41]:
${\cal L}\supset-\frac{1}{4}\frac{1}{\tau_{i}^{2}}\partial_{\mu}a_{i}\partial^{\mu}a_{i}=-\frac{1}{4}e^{-2\sqrt{2}u_{i}}\partial_{\mu}a_{i}\partial^{\mu}a_{i}\;.$
After expanding, we obtain the result
${\cal L}\supset-\frac{1}{4}\left[\partial_{\mu}a_{i}\partial^{\mu}a_{i}-2\sqrt{2}\left(u_{1}\partial_{\mu}a_{1}\partial^{\mu}a_{1}+u_{2}\partial_{\mu}a_{2}\partial^{\mu}a_{2}+u_{3}\partial_{\mu}a_{3}\partial^{\mu}a_{3}\right)\right]\;,$
where the first term is the pure kinetic energy term for axions, while the second corresponds to the interaction terms. Considering the interaction part and expressing the values $u_{1}$,$u_{2}$ and $u_{3}$ in terms of u, v, and t we have,
${\cal L}$ $\supset$ $\frac{1}{\sqrt{6}}t\left(\partial_{\mu}a_{1}\partial^{\mu}a_{1}+\partial_{\mu}a_{2}\partial^{\mu}a_{2}+\partial_{\mu}a_{3}\partial^{\mu}a_{3}\right)$ (4.4)
$+\frac{1}{2}u\left(\partial_{\mu}a_{1}\partial^{\mu}a_{1}-\partial_{\mu}a_{2}\partial^{\mu}a_{2}\right)$
$+\frac{1}{2\sqrt{3}}v\left(\partial_{\mu}a_{1}\partial^{\mu}a_{1}+\partial_{\mu}a_{2}\partial^{\mu}a_{2}-2\partial_{\mu}a_{3}\partial^{\mu}a_{3}\right).$
@@
where $m_{u}$ is the modulus mass. In the large volume scenario, a distinct hierarchy of mass scales is generated. Here, we employ Kähler log corrections in stabilizing the moduli fields, yet the overall spectrum remains consistent with the findings discussed in [42]. The mass eigenstates after diagonalization in units of $M_{p}=1$ can be written as,
@@
where $g_{*}(T_{R}h)$ is the effective degree of freedom at the time of reheating the Universe. The measured values of $N_{\text{eff}}$ require $\Delta N_{\text{eff}}\lesssim 0.95$ at the $95\%$ confidence level, which translates into a bound on the parameters of the model, $a\cdot\lambda\gtrsim 0.1688$. In Fig. 3, the parametric space predicted by the inflationary analysis has $0.86\leq a\leq 1$. This range aligns with the values predicted by $\Delta N_{\text{eff}}\lesssim 0.95$ and is consistent with dark radiation. Within this parametric space, we obtain a tensor-to-scalar ratio values of $r\leq 1.4\times 10^{-4}$, $1\times 10^{15}\lesssim M\lesssim 1.3\times 10^{16}$ GeV, and a soft mass parameter $4.6\times 10^{5}\lesssim M_{s_{c}}\lesssim 1\times 10^{11}$ GeV.
@@ 5 Conclusions
Appendix A Soft Term Potential
In general, the soft term scalar Lagrangian can be written as
$\mathcal{L}\supset m_{\alpha}^{2}C^{\alpha}\overline{C}^{\bar{\alpha}}$ (A.1)
where the soft mass $m_{\alpha}^{2}$ of the field $C$ defined as
$m^{2}_{\alpha}=(m^{2}_{3/2}+V_{0})-F^{\bar{m}}F^{n}\partial_{\bar{m}}\partial_{n}\log\tilde{K}_{\alpha}.$ (A.2)
Here, $F^{\bar{m}}$ is defined as $F^{\bar{m}}=e^{G/2}K^{\bar{m}n}\partial G/\partial n$. Utilizing the superpotential and Kähler metric as given in Eq.(2.4) and Eq.(2.9), we find that the soft mass of the canonically normalized field s is
$m^{2}_{s}\sim(m^{2}_{3/2}+V_{0})-\frac{2}{3{\cal V}_{o}^{2}}W_{o}^{2}+\cdots$ (A.3)
where $V_{0}$ represents the minima of the potential. Hence the soft-term potential is
$\Delta V_{soft}=m^{2}_{S}S^{2}=\left[(m^{2}_{3/2}+V_{0})-\frac{2}{3{\cal V}_{o}^{2}}W_{o}^{2}+\cdots\right]S^{2}.$ (A.4)
Taking into account the relation $s^{2}=\frac{(3a\gamma)}{{\cal{V}}_{0}^{2/3}}S^{2}$ between canonically normalized field s and non-canonically normalized field S we have
$\Delta V_{soft}=\left[(m^{2}_{3/2}+V_{0})-\frac{2}{3{\cal V}_{o}^{2}}W_{o}^{2}+\cdots\right]\frac{{\cal{V}}_{0}^{2/3}}{(3a\gamma)}s^{2}.$ (A.5)
If we further define the ratio $y=s/M$, then A.4 takes the form
$\Delta V_{\text{soft}}=\left[(m^{2}_{3/2}+V_{0})-\frac{2}{3{\cal V}_{o}^{2}}W_{o}^{2}+\cdots\right]M^{2}y^{2}{\frac{{\cal V}_{o}^{2/3}}{3a\gamma}}=M_{s_{c}}^{2}y^{2}~{},$ (A.6)
where
$M_{s_{c}}=\left[(m^{2}_{3/2}+V_{0})-\frac{2}{3{\cal V}_{o}^{2}}W_{o}^{2}+\cdots\right]M^{2}{\frac{{\cal V}_{o}^{2/3}}{3a\gamma}}~{}.$
Appendix B Coleman-Weinberg Corrections
@@
$\small V_{\text{eff}}=\frac{\kappa^{2}\alpha\mid\left(M^{2}-\Phi_{1}\Phi_{2}\right)\mid^{2}+2\gamma\kappa^{2}S^{2}(\alpha\Phi_{1}\Phi_{1}^{\dagger}+\beta\Phi_{2}\Phi_{2}^{\dagger})}{3a\alpha\beta\gamma\mathcal{V}^{4/3}}+\frac{3W_{0}^{2}(2\eta_{0}\log(\mathcal{V})-8\eta_{0}+\xi_{0})}{2\mathcal{V}^{3}}+\frac{d}{\mathcal{V}^{2}},$ (B.1)
@@
$M_{S}^{2}=\begin{pmatrix}\frac{\kappa^{2}S^{2}}{6a\beta\mathcal{V}^{4/3}}&0&-\frac{\kappa^{2}M^{2}}{3a\gamma\mathcal{V}^{4/3}}&0&0&0&0\\ 0&\frac{\kappa^{2}S^{2}}{6a\beta\mathcal{V}^{4/3}}&0&\frac{\kappa^{2}M^{2}}{3a\gamma\mathcal{V}^{4/3}}&0&0&0\\ -\frac{\kappa^{2}M^{2}}{3a\gamma\mathcal{V}^{4/3}}&0&\frac{\kappa^{2}S^{2}}{6a\alpha\mathcal{V}^{4/3}}&0&0&0&0\\ 0&\frac{\kappa^{2}M^{2}}{3a\gamma\mathcal{V}^{4/3}}&0&\frac{\kappa^{2}S^{2}}{6a\alpha\mathcal{V}^{4/3}}&0&0&0\\ 0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0\\ 0&0&0&0&0&0&\frac{28\kappa^{2}M^{4}}{27a\gamma\mathcal{V}^{\frac{10}{3}}}+\cdots\\ \end{pmatrix}$ (B.2)
@@
$\small M_{S}^{2}=\left[0,0,\frac{\kappa^{2}M^{2}\left(y^{2}-\frac{2\alpha}{\gamma}\right)}{6a\mathcal{V}^{4/3}\alpha},\frac{\kappa^{2}M^{2}\left(y^{2}-\frac{2\alpha}{\gamma}\right)}{6a\mathcal{V}^{4/3}\alpha},\frac{\kappa^{2}M^{2}\left(y^{2}+\frac{2\alpha}{\gamma}\right)}{6a\mathcal{V}^{4/3}\alpha},\frac{\kappa^{2}M^{2}\left(y^{2}+\frac{2\alpha}{\gamma}\right)}{6a\mathcal{V}^{4/3}\alpha},\quad\frac{28\kappa^{2}M^{4}}{27a\gamma\mathcal{V}^{10/3}}+\cdots\right]$ (B.3)
@@
$M_{F_{ij}}=\begin{pmatrix}-\frac{6a\gamma\kappa M^{2}S}{\mathcal{V}^{2/3}}+\frac{9a^{2}\gamma^{2}S^{2}}{\mathcal{V}^{4/3}}+\cdots&0&0&\frac{2kM^{2}}{\mathcal{V}}+\frac{2a\gamma S\left(kM^{2}S-4\right)}{\mathcal{V}^{5/3}}+\cdots\\ 0&0&\kappa S&0\\ 0&\kappa S&0&0\\ \frac{2\kappa M^{2}}{\mathcal{V}}+\frac{2a\gamma S\left(\kappa M^{2}S-4\right)}{\mathcal{V}^{5/3}}+\cdots&0&0&\frac{6}{\mathcal{V}^{2}}+\frac{34}{3}a\gamma S^{2}\left(\frac{1}{\mathcal{V}}\right)^{8/3}+\cdots\\ \end{pmatrix}.$ (B.5)
@@
$M_{F_{ij}}^{2}=\left[\kappa^{2}M^{2}y^{2},\kappa^{2}M^{2}y^{2},\frac{36a^{2}\kappa^{2}M^{6}y^{2}\gamma^{2}}{\mathcal{V}^{4/3}}+\cdots,\frac{36a^{2}\kappa^{2}M^{6}y^{2}\gamma^{2}}{\mathcal{V}^{4/3}}+\cdots\right].$ (B.6)
The fermionic mass $\frac{36a^{2}\kappa^{2}M^{6}y^{2}\gamma^{2}}{\mathcal{V}^{4/3}}$ and the bosonic mass $\,\frac{28\kappa^{2}M^{4}}{27a\gamma\mathcal{V}^{10/3}}$ are small compared to other masses, so their contribution to the Coleman-Weinberg potential is suppressed. Additionally, we focus on the region where $(y^{2}\gg 2\alpha/\gamma)$. The relation between canonically and non-canonically normalised field is $s=S\sqrt{3a\gamma}/V_{0}^{1/3}$. The effective contribution of the one-loop radiative corrections can be calculated using the Coleman-Weinberg formula
$\Delta V_{\text{1-loop}}=\frac{1}{64\pi^{2}}\left[M_{S}^{4}\log\left(\frac{M_{S}^{2}}{Q^{2}}\right)-2M_{F}^{4}\log\left(\frac{M_{F}^{2}}{Q^{2}}\right)-\frac{3}{2}\left(M_{S}^{4}-2M_{F}^{4}\right)\right]$ (B.7)
$\Delta V_{\text{1-loop}}=\frac{\kappa^{4}M^{4}y^{4}}{144\pi^{2}a^{2}\gamma^{2}{\cal V}_{o}^{4/3}}\left[F(y)-\left(\frac{1}{54a^{2}\alpha^{2}\gamma^{2}}-\frac{3}{2}{\cal V}_{o}^{8/3}\right)\right],$ (B.8)
@@
$F(y)=\frac{1}{81a^{2}\alpha^{2}\gamma^{2}}\log\left(\frac{\kappa^{2}M^{2}y^{2}}{27a^{2}\alpha\gamma^{2}Q^{2}{\cal V}_{o}^{2/3}}\right)-{\cal V}_{o}^{8/3}\log\left(\frac{\kappa^{2}M^{2}y^{2}{\cal V}_{o}^{2/3}}{3a\gamma Q^{2}}\right)~{}.$ (B.9)